OpenStudy (anonymous):

is 64c^15-d^27 a difference of cubes? @mathmate

OpenStudy (anonymous):

quickly help me please :(

OpenStudy (mathmate):

@zahraaalfa First you decide if \(64c^{15}\) is a perfect cube by checking if the exponents are divisible by 3. Then check using the same method if \(d^{27}\) is a perfect cube. If they both are, then they are the difference of cubes. Example: \(8x^3-27y^3z^6=2^3x^3-3^3y^3(z^2)^3\) so it is a difference of two cubes, we can continue to factorize: \(=(2x)^3-(3yz^2)^3=(2x-3yz^2)(4x^2+6xyz^2+9y^2z^4)\)